OpenStudy (anonymous):

Decide whether you would use the graphing, substitution, or elimination method to solve the following system of equations. Explain, in complete sentences, why you chose that method. Part 2: Solve the following system of equations and show all of your work. 2x + 5y = −3 −15x − 5y = −10

OpenStudy (anonymous):

elimination because it would be easy to eliminate in that type of quadratic equation

OpenStudy (anonymous):

What about part 2? Can you answer it?

OpenStudy (anonymous):

yes hold on give me a second

OpenStudy (anonymous):

y = -5/6 i think

OpenStudy (anonymous):

no wait

OpenStudy (anonymous):

y = -1

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

How did you get y =-1 ?

OpenStudy (anonymous):

do you know how to eliminate? because all I did was add the first equation to the second, found X, then after finding X you plug it in the equation to find Y.

OpenStudy (anonymous):

Would you mind showing me?

OpenStudy (anonymous):

ok give me like a second.

OpenStudy (anonymous):

-13x=-13

OpenStudy (anonymous):

because I added both equations

OpenStudy (anonymous):

and then u simplify. so X=1 because i divided -13 by -13

OpenStudy (anonymous):

What do you mean you added both equations. i dont understand how you got -13

OpenStudy (anonymous):

then u plug it in to any of the equations. 2(1)=5y=-3 5y= -5 y= -1

OpenStudy (anonymous):

ok so I added both equations together.

OpenStudy (anonymous):

2x +5y = - 3 -15x - 5y = -10

OpenStudy (anonymous):

i add -3 and -10, add 2x and -15x, and add 5y and -5y

OpenStudy (anonymous):

and 5y - 5y=0 so now you are left with -13x =-13

OpenStudy (anonymous):

Ohhhhhhhhhhh Now i get it

OpenStudy (anonymous):

I appreciate the help and breakdown

OpenStudy (anonymous):

yep. thats elimination. but that was an easy one.

OpenStudy (anonymous):



OpenStudy (anonymous):

whenever you use elimination, you always want to get either X or Y by themselves. this means you might have to multiply the first equation by a certain number to get either X or Y equal to 0 when you add them. I suggest going to your teacher for help because its very complicated